# Singular-degenerate parabolic systems with the conormal boundary condition on the upper half space

### v1

@@ 5. Equations with partially BMO coefficients in weighted spaces
For the proof, we use the strategy from [13]. We first show the a-priori estimate (2.5). Assume for a moment that $B^{i}=\hat{B}^{i}=C=0$.
First, from the reverse Hölder property of Muckenhoupt weights, there exist sufficiently small $q_{0},\sigma\in(1,\infty)$ depending only on $p,q,d,\alpha$, and $K$, such that $q_{0}\sigma<\min(p,q)$ and
$w_{0}\in A_{q/(q_{0}\sigma)}(\mathbb{R}),\quad w_{1}\in A_{p/(q_{0}\sigma)}(\mathbb{R}^{d}_{+},\mu),$
@@
$L_{q,p,\text{loc}}(\overline{\mathbb{R}^{d+1}_{+}},wd\mu)\subset L_{q_{0}\sigma,\text{loc}}(\overline{\mathbb{R}^{d+1}_{+}},\mu).$
@@
Moreover, $w_{1}]_{A_{p/(q_{0}\sigma)}(\mathbb{R}^{d}_{+},\mu)}\leq\tilde{K}=\tilde{K}(d,\alpha,p,K)$ and $w_{0}]_{A_{q/(q_{0}\sigma)}(\mathbb{R})}\leq\tilde{K}=\tilde{K}(q,K)$. For the proof, see [26, Theorem 7.2.5].
@@
Let $\delta\in(0,1)$ and $\nu\in[32,\infty)$ be some constants which will be determined later. For the time being, assume that $\operatorname{supp}u\subset(s-(\delta R_{0})^{2},s)\times\mathbb{R}^{d}_{+}$ for some $s\in\mathbb{R}$. Fix any $Q^{m}\in\mathbb{C}_{m}^{+}$. We claim that there exists $U^{C}$ on $Q^{m}$ such that $N^{-1}_{d,n,\kappa}U\leq U^{C}\leq N_{d,n,\kappa}U$ on $Q^{m}$ and for any $z\in Q^{m}$, it holds that
@@
$\begin{gathered}\fint_{Q^{m}}|U^{C}-(U^{C})_{Q^{m}}|d\mu\leq N\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}(\mathcal{M}^{1/{q_{0}}}(|F|^{q_{0}})(z)+\mathcal{M}^{1/{q_{0}}}(|f|^{q_{0}})(z))\\ +N(\nu^{-1}+\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}\gamma_{0}^{\frac{1}{{q_{0}}\sigma^{\prime}}}+\delta^{2-\frac{2}{q_{0}\sigma}}\nu^{2-\frac{2}{q_{0}\sigma}})\mathcal{M}^{1/({q_{0}}\sigma)}(|U|^{{q_{0}}\sigma})(z).\end{gathered}$ (5.3)
@@
Indeed, take any $z_{0}\in\overline{\Omega_{T}}$ and $r>0$ such that $Q^{m}\subset Q_{r}^{+}(z_{0})$ and $r/2^{-m}\in[N_{d}^{-1},N_{d}]$. If $r\in(0,\frac{R_{0}}{\nu}]$, then (5.3) follows by Proposition 5.1 (we take $U^{C}=|U^{Q}|$, where $U^{Q}$ is from Proposition 5.1). If $r>\frac{R_{0}}{\nu}$, we take $U^{C}=U$ and (5.3) follows by Lemma 5.2.
@@
By applying the generalized weighted Fefferman-Stein inequality (Theorem 5.3) and the weighted Hardy-Littlewood inequality (Theorem 4.13), we get
@@
$\begin{gathered}\|U\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}(\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+\|f\|_{L_{q,p}(\Omega_{T},wd\mu)})\\ +N(\nu^{-1}+\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}\gamma_{0}^{\frac{1}{{q_{0}}\sigma^{\prime}}}+\delta^{2-\frac{2}{q_{0}\sigma}}\nu^{2-\frac{2}{q_{0}\sigma}})\|U\|_{L_{q,p}(\Omega_{T},wd\mu)}.\end{gathered}$
@@
Taking $\nu=\max(32,4N)$, $\delta$ with $(\delta\nu)^{2-\frac{2}{q_{0}\sigma}}=1/(8N)$, and sufficiently small $\gamma_{0}$ with $\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}\gamma_{0}^{\frac{1}{{q_{0}}\sigma^{\prime}}}\leq 1/(8N)$, we arrive to (2.5). Note that so far we did not impose any restriction on $\lambda$ other than $\lambda\geq 0$.
@@
To remove the assumption $\operatorname{supp}u\subset(s-(\delta R_{0})^{2},s)\times\mathbb{R}^{d}_{+}$, one can use a standard partition of unity in time argument. Precisely, fix a cut-off function $\xi=\xi(t)\in C_{0}^{\infty}(-(R_{0}\delta)^{2},0)$ with $\int_{\mathbb{R}}\xi(s)^{q}ds=1$ and $\int_{\mathbb{R}}|\xi^{\prime}(s)|^{q}ds\leq\frac{N}{(R_{0}\delta)^{2q}}$. For each $s\in\mathbb{R}_{T}$, apply estimate (2.5) to $u(t,x)\xi(t-s)$, raise it to power $q$ and integrate over $s\in\mathbb{R}_{T}$. Here one needs to take $\lambda\geq\lambda_{0}\geq\frac{N}{(\delta R_{0})^{2}}$ to close the argument. We refer to the proof of Theorem 1.1 in [16] for details. Note that both $N$ and $\delta$ depend on $d,\alpha,p,q,\kappa,$ and $K$.
@@
In the general case, we treat the lower-order terms with $B,\hat{B}$, and $C$ as perturbations. Since $|B|\leq K_{1}x_{d},|\hat{B}|\leq K_{1}x_{d},|C|\leq K_{1}x_{d}^{2}$, the terms $D_{i}(x_{d}^{\alpha-1}B^{i}u)$ and $x_{d}^{\alpha-1}\hat{B}^{i}D_{i}u+x_{d}^{\alpha-2}Cu$ can be combined with respectively $D_{i}(x_{d}^{\alpha}F^{i})$ and $\sqrt{\lambda}x_{d}^{\alpha}f$ in the right hand side of the equation. Taking $\lambda_{0}\geq NK_{1}(1+K_{1})$, estimate (2.5) holds in full generality.
@@
The existence of solutions can be proved by adapting the argument in [13, Section 8]. We give a sketch of the proof. It suffices to consider $T=\infty$. By the method of continuity, it suffices to consider $A^{ij}=\delta^{ij}I_{n\times n}$ and $C_{0}=I_{n\times n}$. We keep $a_{0}(x_{d})$ unchanged as the definitions of the spaces $\mathbb{H}^{-1}_{q,p}$ and $\mathcal{H}^{1}_{q,p}$ depend on $a_{0}(x_{d})$. Again from the reverse Hölder property of Muckenhoupt weights, see [26, Theorem 7.2.5], there exists large enough $p_{1}\in(\max(p,q),\infty)$ such that
@@
Take sequences $F_{k},f_{k}\in C_{0}^{\infty}(\mathbb{R}^{d+1}_{+})$ such that $F_{k}\to F$ and $f_{k}\to f$ in $L_{q,p}(\mathbb{R}^{d+1}_{+},wd\mu)$. By Theorem 4.1, there exist solutions $u_{k}\in\mathcal{H}^{1}_{p_{1}}(\mathbb{R}^{d+1}_{+},\mu)$ with $F_{k},f_{k}$ in place of $F,f$. By using the doubling property of measures $w_{0}^{\frac{p_{1}}{p_{1}-q}}(t)dt$ and $w_{1}^{\frac{p_{1}}{p_{1}-p}}(x)d\mu(x)$, an iteration argument similar to the one used in the proof of Lemma 4.11 shows that $u_{k}\in\mathcal{H}^{1}_{q,p}(\mathbb{R}^{d+1}_{+},wd\mu)$. See also [13, Section 8] for details on the iteration argument in the mixed-norm weighted setting. By the a-priori estimate (2.5), the sequence $u_{k}$ is Cauchy in $\mathcal{H}^{1}_{q,p}(\mathbb{R}^{d+1}_{+},wd\mu)$, and therefore converges to the solution $u\in\mathcal{H}^{1}_{q,p}(\mathbb{R}^{d+1}_{+},wd\mu)$ with $F,f$. ∎
@@ 6. Equations with singular lower order terms
Lemma 6.1 and Lemma 6.2 are special cases of [40, Theorem 2] and [40, Theorem 1], respectively.
@@
$D_{i}(x_{d}^{\alpha-1}B^{i}u)=D_{i}(x_{d}^{\alpha}(B^{i}/x_{d})I_{|B|\leq K_{1}x_{d}}u)+D_{i}(x_{d}^{\alpha}B^{i}I_{K_{1}x_{d}<|B|\leq\varepsilon}u/x_{d})={\rm I}+{\rm II},$
$x_{d}^{\alpha-2}Cu=x_{d}^{\alpha}(C/x_{d}^{2})I_{|C|\leq K_{1}x_{d}^{2}}u+x_{d}^{\alpha}(C/x_{d})I_{K_{1}x_{d}^{2}<|C|\leq K_{1}x_{d}}u/x_{d}={\rm I}^{\prime}+{\rm II}^{\prime}.$
By combining the terms $\rm II$} and $\rm II$$^{\prime}$ with respectively $D_{i}(x_{d}^{\alpha}F^{i})$ and $\sqrt{\lambda}x_{d}^{\alpha}f$ in the right hand side of the equation, using (2.5) and (6.1), and taking sufficiently small $\varepsilon$ and sufficiently large $\lambda_{0}$, we arrive to (2.6). ∎
@@
$\|G_{0}\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N\|F_{0}\|_{L_{q,p}(\Omega_{T},wd\mu)}.$ (6.2)
Then for any $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d})$, we have
@@
Hence, both sides of (2.3), viewed as functionals on $C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d})$, are indeed elements of $\mathbb{H}^{-1}_{q,p}(\Omega_{T},wd\mu)$. This justifies the application of the method of continuity between $\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu)$ and $\mathbb{H}^{-1}_{q,p}(\Omega_{T},wd\mu)$. It remains to prove the a-priori estimate (2.5) as the existence of solutions would then follow from the method of continuity and the existence from Theorem 2.2 (i). Note that if $|B|\leq K_{1}x_{d},|\hat{B}|\leq K_{1}x_{d}$, and $|C|\leq K_{1}x_{d}^{2}$, estimate (2.5) follows from Theorem 2.2 (i), (6.3), and (6.2). Otherwise, we split
$x_{d}^{\alpha-1}\hat{B}^{i}D_{i}u=x_{d}^{\alpha}(\hat{B}^{i}/x_{d})I_{|\hat{B}|\leq K_{1}x_{d}}D_{i}u+x_{d}^{\alpha-1}\hat{B}^{i}I_{K_{1}x_{d}<|\hat{B}|\leq\varepsilon}D_{i}u={\rm I}+{\rm II},$
$x_{d}^{\alpha-2}Cu=x_{d}^{\alpha}(C/x_{d}^{2})I_{|C|\leq K_{1}x_{d}^{2}}u+x_{d}^{\alpha-1}(C/x_{d})I_{K_{1}x_{d}^{2}<|C|\leq K_{1}x_{d}}u={\rm I}^{\prime}+{\rm II}^{\prime}.$
By combining the terms $\rm II$} and $\rm II$$^{\prime}$ with $x_{d}^{\alpha-1}F_{0}$ on the right-hand side of the equation and taking sufficiently small $\varepsilon$ and sufficiently large $\lambda_{0}$, estimate (2.5) holds in full generality.
(ii) Again, we only need to prove the a-priori estimate (2.6) as the existence of solutions would then follow from the method of continuity and the existence in Part (i). As $w_{3}\in X_{p}(\mu)$, (6.1) holds. Note that if $|C|\leq\max(K_{1}x_{d},K_{1}x_{d}^{2})$, (2.6) follows by combining Part (i), Theorem 2.2 (ii), and (6.1). Otherwise, we split
$x_{d}^{\alpha-2}Cu=x_{d}^{\alpha-2}CI_{|C|\leq\max(K_{1}x_{d},K_{1}x_{d}^{2})}u+x_{d}^{\alpha-1}CI_{\max(K_{1}x_{d},K_{1}x_{d}^{2})<|C|\leq\varepsilon}u/x_{d}={\rm I}+{\rm II},$
absorb the $\rm II$} term into $x_{d}^{\alpha-1}F_{0}$, and take sufficiently small $\varepsilon$. ∎
We finish the paper with the proof of Remark 2.4 (6). We first present the following lemma that shows that $C_{0}^{\infty}(\mathbb{R}^{d}_{+})$ is dense in $W^{1}_{p}(\mathbb{R}^{d}_{+},wd\mu)$ if $w_{3}\in X_{p}(\mu)$.

### v2

@@ 5. Equations with partially BMO coefficients in weighted spaces
For the proof, we use the strategy from [14]. We first show the a-priori estimate (2.5). Let $F^{i},f\in L_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$ and $u\in\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$ satisfy (1.1)-(1.2). Assume for a moment that $B^{i}=\hat{B}^{i}=C=0$.
First, from the reverse Hölder property of Muckenhoupt weights, there exists $p_{0}=p_{0}(d,\alpha,p,q,K)\in(1,\min(p,q))$ sufficiently close to 1 such that
$w_{0}\in A_{q/p_{0}}(\mathbb{R}),\quad w_{1}\in A_{p/p_{0}}(\mathbb{R}^{d}_{+},\mu),$ (5.3)
@@
$L_{q,p,\text{loc}}(\overline{\mathbb{R}^{d+1}_{+}},wd\mu)\subset L_{p_{0},\text{loc}}(\overline{\mathbb{R}^{d+1}_{+}},\mu).$
@@
Moreover, $w_{1}]_{A_{p/p_{0}}(\mathbb{R}^{d}_{+},\mu)}\leq\tilde{K}=\tilde{K}(d,\alpha,p,K)$ and $w_{0}]_{A_{q/p_{0}}(\mathbb{R})}\leq\tilde{K}=\tilde{K}(q,K)$. For the proof, see [29, Theorem 7.2.5]. Fix any $q_{0}>1$ and $\sigma>1$ with $q_{0}\sigma=p_{0}$; for instance, take $q_{0}=\sigma=p_{0}^{1/2}$.
@@
We first consider the case $0<R_{0}<\infty$. Let $\delta>0$ and $\nu\in[32,\infty)$ be some constants which will be determined later. For the time being, assume that $\operatorname{supp}u\subset(s-(\delta R_{0})^{2},s)\times\mathbb{R}^{d}_{+}$ for some $s\in\mathbb{R}$. Fix any $Q^{m}\in\mathbb{C}_{m}^{+}$. We claim that there exists $U^{C}$ on $Q^{m}$ such that $N^{-1}_{\kappa}U\leq U^{C}\leq N_{\kappa}U$ on $Q^{m}$ and for any $z\in Q^{m}$ it holds that
@@
$\begin{gathered}\fint_{Q^{m}}|U^{C}-(U^{C})_{Q^{m}}|d\mu\leq N\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}(\mathcal{M}^{1/{q_{0}}}(|F|^{q_{0}})(z)+\mathcal{M}^{1/{q_{0}}}(|f|^{q_{0}})(z))\\ +N(\nu^{-1}+\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}\gamma_{0}^{\frac{1}{{q_{0}}\sigma^{\prime}}}+(\delta\nu)^{2-\frac{2}{q_{0}\sigma}})\mathcal{M}^{1/({q_{0}}\sigma)}(|U|^{{q_{0}}\sigma})(z).\end{gathered}$ (5.4)
@@
Let $z_{0}\in\overline{Q^{m}}$ be the top-right corner of $\overline{Q^{m}}$ and let $r=2^{\frac{d-1}{2}}2^{-m}$ so that $Q^{m}\subset Q_{r}^{+}(z_{0})$. If $r\in(0,\frac{R_{0}}{\nu}]$, then (5.4) follows by Proposition 5.1 (we take $U^{C}=U^{Q}$, where $U^{Q}$ is from Proposition 5.1). If $r>\frac{R_{0}}{\nu}$, we take $U^{C}=U$ and (5.4) follows by Lemma 5.2.
@@
Using (5.3), we apply the generalized weighted Fefferman-Stein inequality (Theorem 5.3) and the weighted Hardy-Littlewood inequality (Theorem 4.13), and get
@@
$\begin{gathered}\|U\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq N\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}(\|F\|_{L_{q,p}(\Omega_{T},wd\mu)}+\|f\|_{L_{q,p}(\Omega_{T},wd\mu)})\\ +N(\nu^{-1}+\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}\gamma_{0}^{\frac{1}{{q_{0}}\sigma^{\prime}}}+(\delta\nu)^{2-\frac{2}{q_{0}\sigma}})\|U\|_{L_{q,p}(\Omega_{T},wd\mu)}.\end{gathered}$
@@
Fixing $\nu=\max(32,4N)$, $\gamma_{0}>0$ with $\nu^{\frac{d+2+\alpha_{+}}{{q_{0}}}}\gamma_{0}^{\frac{1}{{q_{0}}\sigma^{\prime}}}=1/(8N)$, and $\delta>0$ with $(\delta\nu)^{2-\frac{2}{q_{0}\sigma}}=1/(8N)$, we arrive to (2.5). Note that so far we did not impose any restriction on $\lambda$ other than $\lambda\geq 0$.
@@
To remove the assumption $\operatorname{supp}u\subset(s-(\delta R_{0})^{2},s)\times\mathbb{R}^{d}_{+}$, one can use a standard partition of unity in time argument. Precisely, fix a cut-off function $\xi=\xi(t)\in C_{0}^{\infty}((-(R_{0}\delta)^{2},0))$ with $\int_{\mathbb{R}}\xi(s)^{q}ds=1$ and $\int_{\mathbb{R}}|\xi^{\prime}(s)|^{q}ds\leq\frac{N}{(R_{0}\delta)^{2q}}$. For each $s\in\mathbb{R}_{T}$, apply estimate (2.5) to $u(t,x)\xi(t-s)$, raise it to power $q$ and integrate over $s\in\mathbb{R}_{T}$. Here one needs $\lambda\geq\frac{N}{(\delta R_{0})^{2}}$ to close the argument. We set $\lambda_{0}:=\frac{N}{\delta^{2}}$. Note that $\delta$, all constants $N$, and $\lambda_{0}$ depend on $d,\alpha,p,q,\kappa,$ and $K$.
If $R_{0}=\infty$, assumption $\operatorname{supp}u\subset(s-(\delta R_{0})^{2},s)\times\mathbb{R}^{d}_{+}$ is not needed and estimate (5.4) holds without the term $(\delta\nu)^{2-\frac{2}{q_{0}\sigma}}$. Hence, if $R_{0}=\infty$, estimate (2.5) holds for any $\lambda\geq 0$.
@@
In the general case, we treat the lower-order terms with $B,\hat{B}$, and $C$ as perturbations. Let $N_{0}$ be the constant in estimate (2.5) for $B^{i}=\hat{B}^{i}=C=0$ and let $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$, where the constant $M>0$ will be determined below. Since $|B|\leq K_{1}x_{d},|\hat{B}|\leq K_{1}x_{d},|C|\leq K_{1}^{2}x_{d}^{2}$, the terms $D_{i}(x_{d}^{\alpha-1}B^{i}u)$ and $x_{d}^{\alpha-1}\hat{B}^{i}D_{i}u+x_{d}^{\alpha-2}Cu$ can be combined with respectively $D_{i}(x_{d}^{\alpha}F^{i})$ and $\sqrt{\lambda}x_{d}^{\alpha}f$ on the right-hand side of the equation. This gives ($\|\cdot\|=\|\cdot\|_{L_{q,p}(\Omega_{T},wd\mu)}$)
$\begin{gathered}\|Du\|+\lambda^{1/2}\|u\|\leq N_{0}\|F\|+N_{0}\|f\|+N_{0}K_{1}\|u\|+\lambda^{-1/2}N_{0}K_{1}\|Du\|+\lambda^{-1/2}N_{0}K_{1}^{2}\|u\|\\ \leq N_{0}\|F\|+N_{0}\|f\|+N_{0}M^{-1/2}\|Du\|+N_{0}K_{1}(M^{-1/2}+1)\|u\|.\end{gathered}$
Taking $M$ sufficiently large, for example $M=16N_{0}^{2}+1$, and $N=2N_{0}$, gives estimate (2.5) in full generality.
@@
The existence of solutions can be proved by adapting the argument in [14, Section 8]. We give a sketch of the proof. It suffices to consider $T=\infty$. By the method of continuity, it suffices to consider $A^{ij}=\delta^{ij}I_{n\times n}$ and $C_{0}=I_{n\times n}$. We keep $a_{0}(x_{d})$ unchanged as the definition of the space $\mathcal{H}^{1}_{q,p}$ depends on $a_{0}(x_{d})$. Again from the reverse Hölder property of Muckenhoupt weights, see [29, Theorem 7.2.5], there exists large enough $p_{1}\in(\max(p,q),\infty)$ such that
@@
Take sequences $F_{k},f_{k}\in C_{0}^{\infty}(\mathbb{R}^{d+1}_{+})$ such that $F_{k}\to F$ and $f_{k}\to f$ in $L_{q,p}(\mathbb{R}^{d+1}_{+},wd\mu)$. By Theorem 4.1, there exist solutions $u_{k}\in\mathcal{H}^{1}_{p_{1}}(\mathbb{R}^{d+1}_{+},\mu;\mathbb{R}^{n})$ with $F_{k},f_{k}$ in place of $F,f$. By using the doubling property of measures $w_{0}^{\frac{p_{1}}{p_{1}-q}}(t)dt$ and $w_{1}^{\frac{p_{1}}{p_{1}-p}}(x)d\mu(x)$, an iteration argument similar to the one used in the proof of Lemma 4.11 shows that $u_{k}\in\mathcal{H}^{1}_{q,p}(\mathbb{R}^{d+1}_{+},wd\mu;\mathbb{R}^{n})$. See also [14, Section 8] for details on the iteration argument in the mixed-norm weighted setting. By the a-priori estimate (2.5), the sequence $u_{k}$ is Cauchy in $\mathcal{H}^{1}_{q,p}(\mathbb{R}^{d+1}_{+},wd\mu;\mathbb{R}^{n})$, and therefore converges to the solution $u\in\mathcal{H}^{1}_{q,p}(\mathbb{R}^{d+1}_{+},wd\mu;\mathbb{R}^{n})$ with $F,f$. ∎
@@ 6. Equations with singular lower order terms
Lemma 6.1 and Lemma 6.2 are special cases of [47, Theorem 2] and [47, Theorem 1], respectively.
@@
$D_{i}(x_{d}^{\alpha-1}B^{i}u)=D_{i}(x_{d}^{\alpha}(B^{i}/x_{d})I_{|B|\leq K_{1}x_{d}}u)+D_{i}(x_{d}^{\alpha}B^{i}I_{K_{1}x_{d}<|B|\leq\varepsilon K_{2}^{-1}}u/x_{d})={\rm I}+{\rm II},$
$x_{d}^{\alpha-2}Cu=x_{d}^{\alpha}(C/x_{d}^{2})I_{|C|\leq K_{1}^{2}x_{d}^{2}}u+x_{d}^{\alpha}(C/x_{d})I_{K_{1}^{2}x_{d}^{2}<|C|\leq K_{1}K_{2}^{-1}x_{d}}u/x_{d}={\rm I}^{\prime}+{\rm II}^{\prime}.$
Let $N_{1}>0$ be the constant in estimate (2.5) and let $M_{1}$ be the constant $M$ from Theorem 2.2 (i). Assume that $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$, where the constant $M\geq M_{1}$ will be determined below. By combining the terms $\rm II$} and $\rm II$$^{\prime}$ with respectively $D_{i}(x_{d}^{\alpha}F^{i})$ and $\sqrt{\lambda}x_{d}^{\alpha}f$ in the right hand side of the equation, using (2.5) and (6.1), we get ($\|\cdot\|=\|\cdot\|_{L_{q,p}(\Omega_{T},wd\mu)}$)
$\begin{gathered}\|Du\|+\sqrt{\lambda}\|u\|\leq N_{1}\|F\|+N_{1}\|f\|+N_{1}(\varepsilon K_{2}^{-1}+\lambda^{-1/2}K_{1}K_{2}^{-1})\|u/x_{d}\|\\ \leq N_{1}\|F\|+N_{1}\|f\|+2N_{1}(\varepsilon+M^{-1/2})\|D_{d}u\|.\end{gathered}$
Taking $\varepsilon=1/(8N_{1})$, $M=\max(64N_{1}^{2},M_{1})$, and using (6.1), we arrive to (2.6) with $N=6N_{1}$. ∎
@@
$\|G_{0}\|_{L_{q,p}(\Omega_{T},wd\mu)}\leq 2[w_{3}]_{Y_{p}(\mu)}\|F_{0}\|_{L_{q,p}(\Omega_{T},wd\mu)}.$ (6.2)
Then for any $\varphi\in C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d};\mathbb{R}^{n})$, we have
@@
Hence, both sides of (2.3), viewed as functionals on $C_{0}^{\infty}((-\infty,T)\times\mathbb{R}^{d};\mathbb{R}^{n})$, are indeed elements of $\mathbb{H}^{-1}_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$. This justifies the application of the method of continuity between $\mathcal{H}^{1}_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$ and $\mathbb{H}^{-1}_{q,p}(\Omega_{T},wd\mu;\mathbb{R}^{n})$. It suffices to prove the a-priori estimate (2.7) as the existence of solutions would then follow from the method of continuity and the existence from Theorem 2.2 (i).
Let $N_{1}>0$ be the constant in estimate (2.5) and let $M_{1}$ be the constant $M$ from Theorem 2.2 (i). Assume that $\lambda\geq\max(\lambda_{0}R_{0}^{-2},MK_{1}^{2})$, where the constant $M\geq M_{1}$ will be determined below. We split
$x_{d}^{\alpha-1}\hat{B}^{i}D_{i}u=x_{d}^{\alpha}(\hat{B}^{i}/x_{d})I_{|\hat{B}|\leq K_{1}x_{d}}D_{i}u+x_{d}^{\alpha-1}\hat{B}^{i}I_{K_{1}x_{d}<|\hat{B}|\leq\varepsilon K_{3}^{-1}}D_{i}u={\rm I}+{\rm II},$
$x_{d}^{\alpha-2}Cu=x_{d}^{\alpha}(C/x_{d}^{2})I_{|C|\leq K_{1}^{2}x_{d}^{2}}u+x_{d}^{\alpha-1}(C/x_{d})I_{K_{1}^{2}x_{d}^{2}<|C|\leq K_{1}K_{3}^{-1}x_{d}}u={\rm I}^{\prime}+{\rm II}^{\prime}.$
By combining the terms $\rm II$} and $\rm II$$^{\prime}$ with $x_{d}^{\alpha-1}F_{0}$ on the right-hand side of the equation, using (6.3) and estimates (6.2), (2.5), we have ($\|\cdot\|=\|\cdot\|_{L_{q,p}(\Omega_{T},wd\mu)}$)
$\begin{gathered}\|Du\|+\sqrt{\lambda}\|u\|\leq N_{1}\|F\|+N_{1}\|f\|+2N_{1}K_{3}(\|F_{0}\|+\varepsilon K_{3}^{-1}\|Du\|+K_{1}K_{3}^{-1}\|u\|).\end{gathered}$
Taking $\varepsilon=1/(4N_{1})$, $M=\max(16N_{1}^{2},M_{1})$, and $N=4N_{1}$, we arrive to estimate (2.7).
(ii) Again, we only need to prove the a-priori estimate (2.8). As $w_{3}\in X_{p}(\mu)\cap Y_{p}(\mu)$, we can derive estimate (2.8) with appropriate constants $\varepsilon,M$, and $N$, by combining Part (i), Theorem 2.2 (ii), and splitting
$x_{d}^{\alpha-2}Cu=x_{d}^{\alpha-2}CI_{|C|\leq h(x_{d})}u+x_{d}^{\alpha-1}CI_{h(x_{d})<|C|\leq\varepsilon(K_{2}K_{3})^{-1}}u/x_{d}={\rm I}+{\rm II},$
where $h(x_{d}):=\max(K_{1}K_{2}^{-1}x_{d},K_{1}K_{3}^{-1}x_{d},K_{1}^{2}x_{d}^{2})$, and we absorb the $\rm II$} term into $x_{d}^{\alpha-1}F_{0}$. ∎
We finish this Section with the proof of Remark 2.4 (6). We first present the following lemma that shows that $C_{0}^{\infty}(\mathbb{R}^{d}_{+})$ is dense in $W^{1}_{p}(\mathbb{R}^{d}_{+},wd\mu)$ if $w_{3}\in X_{p}(\mu)$.